If the force between the two charges is 60N, how the force if the distance is reduced by third and the two charges remain constant?
1
Expert's answer
2020-10-07T09:00:40-0400

Solution.

"F_1=60N,"

"q_1=q_2=const,"

"r_2=r_1-\dfrac{r_1}{3};"

"F_1=k\sdot\dfrac{q_1\sdot q_2}{r_1^2};"

"F_2=k\sdot\dfrac{q_1\sdot q_2} {(\dfrac{2r}{3})^2};"

"F_2=\dfrac{9F_1}{4};"

"F_2=\dfrac{9\sdot60}{4}=135N;"

Answer: "F_2=135N."